OpenStudy (anonymous):

I-15 + x I = 11 solve for X

OpenStudy (hoblos):

-15 +x = 11 -15 +x =-11 x= 26 x=4

OpenStudy (anonymous):

thanks again, thats the answer i got as well. Its good to see when someone else confirms you did it right!

OpenStudy (hoblos):

my pleasure XD